OpenStudy (anonymous):

WILL AWARD MEDAL Addison currently has an account balance of $1,737.83. She opened the account 11 years ago with a deposit of $1,334.62. If the interest compounds semi-annually, what is the interest rate on the account?

OpenStudy (anonymous):

1.2% 2.4% 4.1% 7.7%

OpenStudy (anonymous):

@triciaal

OpenStudy (triciaal):

do you know how to compound interest? remember semi-annually means 2 times per year (every 6 months)

OpenStudy (anonymous):

its been awhile since ive done it

OpenStudy (triciaal):

same here; need the formula for compound interest. for simple interest the formula is I = PRT/100 where I is the interestin $, P is the starting amount in $, R is the rate as a % and T is the time in years

OpenStudy (triciaal):

using simple interest for this problem 403.21 = 1334.62 *r * 22 r =

OpenStudy (triciaal):

I= P(1 + R)^t let me check the formula, i'll be back

OpenStudy (triciaal):

A =P(1 + r)^t is the correct formula for compound interest where A is the accumulated amount after t years

OpenStudy (triciaal):

plug the numbers and calculate

OpenStudy (triciaal):

r u trying it to confirm a or b ?